OpenStudy (anonymous):

I'm having trouble with convert units of weight and mass. Can you teach me how to do it. My homework have like 5T=___lb what does T stands for in math. ps I'm a 6th grader

OpenStudy (phi):

T stands for what ? (I would guess ton ?)

OpenStudy (anonymous):

I think so I'm not really sure but can you teach me how?

OpenStudy (phi):

do you know how many lbs (pounds) per ton ?

OpenStudy (phi):

to do conversions, you need to know that. Your book must tell you ?

OpenStudy (phi):

Step 1 is to know the "conversion factor", in this case pounds per ton

OpenStudy (anonymous):

On my paper 1 ton (ib) = 16 ounces (oz)

OpenStudy (phi):

1 pound (abbreviated lb) = 16 ounces (oz) look some more .

OpenStudy (anonymous):

Okay thanks

OpenStudy (phi):

Does your paper have T ?

OpenStudy (anonymous):

Oh I see it yes it it

OpenStudy (phi):

what does it say?

OpenStudy (anonymous):

1ton(T)=2,000 what does dg stands for

OpenStudy (phi):

it should say 1 ton (T) = 2000 pounds (lb) where does it say dg ?

OpenStudy (anonymous):

Nvm I see it lol but can u show me an example

OpenStudy (phi):

OK you want to "change" 5 tons into pounds 5T=___lb we will use this idea: anything divided by itself is 1 if we have \[ ton \cdot \frac{pounds}{ton} \] then we can use this idea to say ton/ton is 1, and we are left with pounds \[ \cancel{ton} \cdot \frac{pounds}{\cancel{ton}} = pounds\]

OpenStudy (phi):

That is the first idea. The second idea is the conversion factor can be written as a fraction for example: 1 ton = 2000 pounds we can write this information as \[ \frac{1 \text{ ton}}{2000 \text{ pounds}} \] or \[ \frac{2000 \text{ pounds}}{1 \text{ ton}} \]

OpenStudy (phi):

Because you want to change tons to pounds, you want to use the second way of writing the fraction (so we can use the first idea) we write \[ 5 \text{tons} \cdot \frac{2000 \text{ pounds}}{1 \text{ ton}} \] "cancel" the tons because ton/ton = 1 \[ 5 \cancel{\text{tons}} \cdot \frac{2000 \text{ pounds}}{1 \cancel{\text{ ton}}} \] or just \[ 5 \cdot 2000 \text{ pounds} \] simplify by doing the multiplication 10,000 pounds